# Federated Unlearning: a Perspective of Stability and Fairness

### v1

@@ Appendix F Balancing fairness Unlearning Algorithm Analysis
Algorithm 1 Balancing Fairness Federated Unlearning
0:  $T,\eta,\epsilon,\Lambda$
1:  Initialize $\boldsymbol{\lambda}=\boldsymbol{0}$.
2:  for $t=0,\cdots,T-1$ do
3:     Server broadcasts $\boldsymbol{w}^{t},{\lambda_{i}}$ to each client $i\notin\mathcal{J}$.
4:     for each client $i$ do
4:         # Local Training Start.
5:        $r_{i}=f_{i}(\boldsymbol{w}^{t})-f_{i}(\boldsymbol{\boldsymbol{w}}^{o})$
6:        Update weight $\boldsymbol{w}_{i}$ for $E$ local epochs with $f^{\prime}(\boldsymbol{w})=f_{i}\left(\boldsymbol{w}\right)+{\lambda_{i}}\left(f_{i}\left(\boldsymbol{w}\right)-f_{i}(\boldsymbol{\boldsymbol{w}}^{o})\right)$
7:        $\boldsymbol{w}_{i}^{(t,\tau+1)}=\boldsymbol{w}_{i}^{(t,\tau)}-\eta\nabla f^{\prime}_{i}(\boldsymbol{w}_{i}^{(t,\tau)})$
8:        Send $\boldsymbol{w}^{(t,E)}$ and $r_{i}$ back to the server
8:         # Local Training End.
9:     end for
10:     Server aggregates the weight $\boldsymbol{w}^{t+1}=\sum_{i\notin\mathcal{J}}p_{i}^{\prime}\boldsymbol{w}_{i}^{(t,E)}.$
11:     If $\max_{i}{r}_{i}\leq\epsilon$: Return $\boldsymbol{w}^{t+1}$
12:     Update $\lambda_{i}=\Lambda\frac{\exp\left(r_{i}\right)}{1+\sum_{i\notin\mathcal{J}}\exp\left(r_{i}\right)}$ for $\forall i\notin\mathcal{J}$.
13:  end for
14:  Return $\boldsymbol{w}^{T}$
@@
Appendix G Experiemnts
In this section, we present experiments to validate the proposed FU mechanisms. We examine various scenarios and settings to demonstrate the effectiveness and robustness of our approaches. Our implementation utilizes the open-source FL framework Plato Li et al. (2023) for reproducibility.
This appendix provides additional information complementing the experiments discussed in Section 7. It includes a detailed data distribution in these experiments and elaborates on the different data heterogeneities.
Figure 6: Non-IID Class Distribution for MNIST (Section 7) and CIFAR-10: Each client has four classes.
G.1 Heterogeneity between remaining and unlearned clients
Label Distribution $S_{\text{retrain}}$ $S_{\lambda=1}$ $V_{\lambda=1}$
0.1 8.24 7.69 (-0.55) 0.0
0.4 3.45 2.05 (-1.4) 0.0
0.7 0.26 0.23 (-0.03) 0.06
Table 2 demonstrates a correlation between the data heterogeneity level and the system’s stability after unlearning.
A higher data heterogeneity leads to increased instability after unlearning, as indicated by the higher $S_{\text{retrain}}$ values. Notably, applying our FU mechanism with a stability trade-off parameter ($\lambda=1$) results in improved stability, as shown by the reduced $S_{\lambda=1}$ values, while the difference in FU verification remains negligible. This underscores our proposed approach could balance the trade-off between unlearning verification and global stability under varying degrees of data heterogeneity.
G.2 Heterogeneity among remaining clients.
This section explores the fairness in FU under varying levels of heterogeneity among remaining clients, and data between remaining and unlearned clients is also heterogeneous [^5] According to Lemma 6.8, given homogeneous two groups, the fairness primarily depends on data heterogeneity among remaining clients, as we aim to investigate the impact of unlearned clients, we choose heterogeneous two groups setting..
To extend our analysis, we consider the CIFAR-10 dataset Krizhevsky (2009) with class distribution in Figure 6 and utilize the ResNet18 architecture He et al. (2016). CIFAR-10 exhibits greater complexity compared to MNIST. Figure 7 illustrates the impact of unlearning on fairness among remaining clients, with a fairness parameter $\Lambda=1,\epsilon=20$. In this scenario, clients [1, 3] experienced significant utility loss due to their similar data distribution with unlearned clients [8, 9] (as shown in Figure 6). However, our FU mechanism achieves a lower fairness metric ($Q_{\text{our}}=6.91$) compared to the retraining approach ($Q_{\text{retrain}}=10.23$), indicating more equitable utility changes among remaining clients. The verification metric in this case is $V=0.42$, demonstrating that our mechanism enhances fairness even in more complex and heterogeneous environments.
Figure 7: Unlearning Impact on Fairness: ensuring no client’s performance deviates beyond $20\%$; fairness metric ($Q$) of our FU mechanism ($Q_{\text{our}}=6.91$) is lower than that of retrain $Q_{\text{retrain}}=10.23$

### v4

@@ Appendix F Balancing fairness Unlearning Algorithm Analysis
Appendix G Broader Impact and Limitations
This paper contributes to federated learning (FL) through federated unlearning (FU) techniques, addressing the growing need for user data rights and regulatory compliance. Our emphasis on fairness and stability in the unlearning process paves the way for more comprehensive and responsible FL systems development. Additionally, our detailed exploration of the impacts of data heterogeneity on the unlearning process offers insights into the development of more advanced FU methods. This focus ensures that our approach is not only theoretically sound but also practically applicable in diverse real-world scenarios.
Limitaions: Our theoretical framework assumes a specific optimization formulation for the unlearning process, and the insights derived may not generalize to all FU settings or optimization techniques. Our focus has primarily been on the theoretical impacts. To enhance the generalizability of our findings, future research should investigate algorithm-centric methods through further experimentation across a wider range of scenarios. Additionally, developing more comprehensive frameworks that account for diverse factors in FU could provide a deeper understanding and more robust solutions.